On your way back from Argentina to the U.S., you buy a souvenir for your friend....

  1. On your way back from Argentina to the U.S., you buy a souvenir for your friend. You pay 150 Argentine Pesos for it. How much is your cost in US dollars?
  2. How many Euros can you buy with US$1,100?
  3. Which currency is more valuable in terms of a US dollar: Euro or Kuwaiti Dinar?
  4. How much is one Trinidad and Tobago dollar worth in US dollars?
  5. You have to settle an account payable in 1,500,000 Swiss Francs. What will be the cost of doing so in US dollars? vi. Go to the world-wide-web and generate a list of resources that provide current and historical exchange rate data.

Please note that an exchange rate was not provided, therefore you must google the exchange rate.
Answer #1

(i) Current Peso/$ Spot Rate = 44.71 Peso/$ (Source: Google Finance)

Souvenir Value in Peso = 150

Souvenir Value in $ = 150 / 44.71 = $ 3.355

(ii) Current EUR/$ Spot Rate = 0.89 / $ (Source: Google Finance)

Equivalent EURO value of $ 1100 = 1100 x 0.89 = $ 979

(iii) Current EUR/$ Spot Rate = 0.89 / $ and Current Dinar/$ Spot Rate = 0.3 Dinar / $ (Source: Google Finance)

As more EUR can be bought for 1 $ as compared to Kuwaiti Dinar, the former is more valuable as compared to the latter.

(iv) Current Trinidad Tobago $ / USD = 6.77 / $ (Source: Google Finance)

(v) Swiss Franc Payable = 1500000 Sf and Current Sf/$ Spot Rate = 0.99 Sf/$ (Source: Google Finance)

$ Value of Payable = 1500000 / 0.99 = $ 1515151.515